The rate constant of a chemical reaction is 9.79 s-1 at a temperature of 22.4 °C. If the activation energy is 52.6 kJ mol-1, what is the rate constant (in s-1) at 53.6 °C?
Answer #1

rate constant k1 = 9.79 s-1

temperature T1 = 22.4 oC = 295.55 K

T2 = 53.6 oC = 326.75 K

Activation energy = 52.6 kJ/mol

ln (k2 / k1) = Ea / R [1 / T1 - 1 / T2]

ln (k2 / 9.79) = 52.6 / 8.314 x 10^-3 [1 / 295.55 - 1/326.75]

k2 = 75.59

rate constant = 75.6 s-1